OpenStudy (anonymous):

What value of k makes the statement true? k + (–4.1) = –9.7

OpenStudy (anonymous):

@One098 @tinybookworm @TierraStudy

OpenStudy (one098):

subtract -9.7 and (-4.1)

OpenStudy (anonymous):

k + (-4.1) = -9.7 Bring (-4.1) to the right side: k = -9.7 - (-4.1) = -9.7 + 4.1 = -5.6 So k = -5.6

OpenStudy (anonymous):

Thanks.